### v1

# SMiR: Efficient Synthetic Data Pipeline to Improve Multi-image Reasoning

Abstract
Vision-Language Models (VLMs) have shown strong performance in understanding single images, aided by numerous high-quality instruction datasets. However, multi-image reasoning tasks are still under-explored in the open-source community due to two main challenges: (1) scaling datasets with multiple correlated images and complex reasoning instructions is resource-intensive and maintaining quality is difficult, and (2) there is a lack of robust evaluation benchmarks for multi-image tasks. To address these issues, we introduce SMiR, an efficient synthetic data-generation pipeline for multi-image reasoning, and a high-quality dataset generated using this pipeline. Our pipeline efficiently extracts highly correlated images using multimodal embeddings, combining visual and descriptive information and leverages open-source LLMs to generate quality instructions. Using this pipeline, we generated 160K synthetic training samples, offering a cost-effective alternative to expensive closed-source solutions. Additionally, we present SMiR-Bench, a novel multi-image reasoning evaluation benchmark comprising 200 diverse examples across 7 complex multi-image reasoning tasks. SMiR-Bench is multi-turn and utilizes a VLM judge to evaluate free-form responses, providing a comprehensive assessment of model expressiveness and reasoning capability across modalities. We demonstrate the effectiveness of SMiR dataset by fine-tuning several open-source VLMs and evaluating their performance on SMiR-Bench. Our results show that models trained on our dataset outperform baseline models in multi-image reasoning tasks up to 8% with a much more scalable data pipeline. Our codebase is available at https://github.com/togethercomputer/SMiR.
1 Introduction
Vision-Language Models (VLMs) have shown impressive capabilities in tasks involving single images, particularly open-source models that have benefited from high-quality instruction datasets (Laurençon et al., 2024b; Zhang et al., 2023; Xu et al., 2022). However, in multi-image tasks, such as comparing or analyzing relationships between multiple images, the performance of open-source VLMs (Liu et al., 2024b; a; Li et al., 2024a; Awadalla et al., 2023; Yao et al., 2024; Wang et al., 2023b) significantly lags behind that of their closed-source counterparts like GPT-4 (Achiam et al., 2023), Claude 3.5 Sonnet (Anthropic, 2024a), Claude 3 (Anthropic, 2024b), and Gemini 1.5 (Reid et al., 2024).
While we do not have full visibility into how open-source models are trained, a reasonable assumption for their underperformance in multi-image reasoning tasks is the advantage closed-source models have in terms of training datasets. Closed-source developers likely benefit from access to large, high-quality datasets, which are critical for training advanced models capable of reasoning across multiple images. In contrast, open-source models may face limitations in accessing or constructing datasets that are specifically tailored to these tasks, which impacts their overall performance.
One major challenge lies in collecting and curating large-scale datasets with highly correlated images. Identifying meaningful connections between images requires vast amounts of data and sophisticated algorithms, which most multi-image instruction-tuning datasets lack. For example, MANTIS (Jiang et al., 2024) often pairs unrelated images in the same reasoning task, unintentionally simplifying the challenge. Consider a question like: Compare the behaviors and characteristics shown in these images. How do they demonstrate different aspects of survival and adaptation? Previous methods may match it with a relevant image of a dog alongside completely unrelated images like a drone hovering in the air and shipping containers at a port. While the question naturally applies to analyzing an animal’s behavioral adaptations, its pairing with technological and industrial objects (drones, logistics, etc.) forces superficial comparisons that dilute the biological focus of the inquiry. This makes the task easier, as the unrelated items can only be discussed through loose metaphorical connections to survival and adaptation.
Figure 1: Comparison of top (example of MANTIS dataset sharegpt4v-pt_125, which can have a data sample with animals, but are concatenated together with unrelated images) vs. bottom (SMiR dataset where related images are together) for multi-image reasoning.
In contrast, comparing related images—like a series of different animals in their natural habitats—demands a deeper understanding of predator behavior and ecological adaptation. Here, the task requires a more nuanced comparison of hunting techniques, territorial behaviors, and social structures within the same family of species. MMDU-45K (Liu et al., 2024e) attempts to address this by clustering image captions and category tags, but it overlooks the crucial visual relationships. These shortcomings highlight the need for datasets with truly correlated images to foster complex multi-image reasoning tasks that better reflect real-world analytical challenges.
Second, scaling up the number of highly correlated images presents significant challenges. Existing datasets such as MANTIS, MMDU-45K, and MultiInstruct (Xu et al., 2022) require extensive human curation and annotation, resulting in a labor-intensive and time-consuming process. To minimize human effort, researchers have leveraged GPT-4 family models (Peng et al., 2023; Wang et al., 2023a) to generate synthetic datasets—including MMInstruct (Liu et al., 2024d), Multimodal ArXiv (Li et al., 2024b), MIMIC-IT (Li et al., 2023a), StableLLaVA-Instruct (Li et al., 2023c), and SVIT-Instruct (Zhao et al., 2023). However, this method proves expensive and difficult to scale effectively.
Third, evaluating multi-image reasoning is challenging due to the limitations of traditional assessment methods. Multiple-choice formats, used in previous multi-image reasoning benchmarks (Meng et al., 2024; Fu et al., 2024; Wang et al., 2024a; Yue et al., 2024), fail to capture the nuanced reasoning process and justifications behind a model’s answers. This is problematic because models can achieve similar performance on multiple-choice tasks while having vastly different reasoning capabilities. In contrast, free-form responses allow models to demonstrate their thought processes, providing a more comprehensive evaluation of their true reasoning abilities and distinguishing between models that rely on superficial pattern matching and those with a deeper understanding.
To address these challenges, we propose a synthetic data generation pipeline, SMiR, for multi-image reasoning, along with a human-annotated evaluation benchmark, SMiR-Bench. SMiR generates correlated and challenging multi-image reasoning questions, while SMiR-Bench evaluates models using free-response tasks in complex multi-image scenarios.
Our contributions are as follows:
  • We present SMiR, a pipeline for multi-image instruction tuning. To construct SMiR, we apply existing algorithms in a novel way to collect multiple images with high correlation. Our approach combines cluster sampling for data quality robustness and graph iteration sampling for diversity. Both algorithms leverage multimodal embeddings that integrate image and caption information to effectively group correlated images, enabling challenging multi-image instruction tuning scenarios.
  • We develop a scalable synthetic multimodal data generation pipeline that leverages open-source LLMs such as Meta Llama 3.1 70B Instruct Turbo (Dubey et al., 2024). This approach eliminates the need for expensive closed-source models, reducing costs by up to 50 times and increasing speed by up to 10 times (Kirkovska, 2024), while significantly minimizing human annotation efforts.
  • We present a new multimodal evaluation benchmark that uses multi-turn interactions and free-form responses to assess both final answers and reasoning processes in complex multi-image tasks. The evaluation is conducted pairwise, directly comparing model performances. Using reference answers from GPT-4-Turbo and other open-source models as baselines, we observe up to an 8% improvement in win rate for instruction-tuned models trained on the SMiR dataset.
2 Related Works
Vision Language Models
We focus on instruction tuning Vision-Language Models (VLMs) that utilize a pretrained Large Language Model (LLM) backbone because this approach is cost-effective and more accessible for the open-source community. Since the backbone responsible for language understanding is already trained, the overall training process becomes simpler and requires fewer resources. Our primary task involves aligning the vision encoder—typically architectures like Vision Transformer (ViT) (Dosovitskiy, 2020), SigLIP (Zhai et al., 2023), or CLIP (Radford et al., 2021)—with the LLM backbone. This alignment is facilitated through linear layers that connect the vision encoder to the backbone, enabling the integration of visual and textual information. For instance, BLIP-2 (Li et al., 2023b) uses OPT (Zhang et al., 2022) and FLAN-T5 (Chung et al., 2022) as backbones, MiniGPT-4 (Zhu et al., 2023) utilizes Vicuna (Chiang et al., 2023), and Qwen-2-VL (Wang et al., 2024b) employs Qwen-2-1.5B (Yang et al., 2024) as the language backbone. In this paper, we focus on creating a high-quality multi-image reasoning dataset for instruction tuning instead of large-scale interleaved pretraining datasets like OBELICS (Laurençon et al., 2024a), MINT-1T (Awadalla et al., 2024), and LAION-5B (Schuhmann et al., 2022).
Multi-Image Reasoning Data
Recent advancements in multi-image reasoning instruction tuning datasets include MANTIS (Jiang et al., 2024) and MMDU-45K (Liu et al., 2024e), both aiming to improve reasoning capabilities in VLMs. However, these datasets have limitations in their approaches. MANTIS randomly concatenates single image pairs from LLaVA-665k (Liu et al., 2024a), which often results in uncorrelated images within multi-image scenarios, potentially undermining the complexity of reasoning tasks. MMDU-45K attempts to address this issue by utilizing sentence transformers (Reimers, 2019) with description text and clustering techniques to group related images, but does not consider visual components. The dataset is then further enhanced, assisted by GPT-4 to generate comprehensive answers for the grouped images. Building upon these efforts, SMiR introduces a novel approach that leverages both vital visual and caption information to ensure highly correlated images within multi-image sets with the use of open-source LLMs. These scalable methods leads to the generation of more challenging questions that require deeper analysis and understanding of visual relationships, pushing the boundaries of multi-image reasoning capabilities in VLMs.
Datasets Multimodal Embedding Correlated Images No Human-Annotation Open-Source LLM
Mantis
MMDU
SMiR
Table 1: Comparison of datasets highlighting key characteristics and methodologies.
Multi-Image Reasoning Benchmarks
Recent VLM benchmarks (Chiang et al., 2024; Lin et al., 2024; Liu et al., 2024c) have made strides by incorporating free-response evaluations, marking a significant improvement over traditional multiple-choice formats. However, these benchmarks still lack a comprehensive approach that combines automatic, multi-turn, and pairwise evaluation capabilities. Our benchmark addresses this gap, drawing inspiration from Auto-Hard-Auto v0.1 (Li et al., 2024c). We have adapted and expanded this framework to enable robust multimodal evaluation, providing a more holistic assessment of VLM performance across complex, multi-image reasoning tasks. This approach allows for a deeper analysis of both the final answers and the underlying reasoning processes employed by VLMs in real-world SMiR-Bench scenarios.
Figure 2: Our end-to-end pipeline: from sampling and LLM prompting to conversation generation. The example conversation is based on a sports scenario, demonstrating the pipeline’s ability to generate contextually relevant multi-turn dialogues.
3 SMiR: Synthetic Multi-Image Reasoning Data Pipeline
SMiR pipeline efficiently generates complex multi-image reasoning data through a carefully designed methodology that reflects key design considerations and trade-offs. Starting with a large-scale image-caption dataset $D$ containing $N$ image-caption pairs $(I_{i},C_{i})_{i=1}^{N}$, SMiR first constructs multimodal embeddings for each pair. Then, grouping algorithms identify correlations between these embeddings. Finally, open-source LLMs are used to generate complex question-answer pairs based on the sampled correlations. In the following subsections, we describe our approach’s core components—the construction of multimodal embeddings, grouping algorithms, and synthetic data generation implementation—while examining the critical design decisions behind our sampling algorithms, prompts, and data sources that shaped the final SMiR dataset.
3.1 Multimodal Embedding Construction
While embedding-based approaches have been used before to capture relationships between images and text, relying solely on either image or caption embeddings limits our ability to fully understand the connections between visual content and descriptive text. A more comprehensive approach requires incorporating both contexts simultaneously.
To address this, we developed a method that combines image and caption information into a single multimodal embedding. Our approach utilizes SigLIP or CLIP embeddings for the images, combined with corresponding caption embeddings. We formulated the multimodal embedding as follows:
$E_{multimodal}=E_{image}+c\cdot E_{caption}$ (1)
where $E_{multimodal}$ is the multimodal embedding, $E_{image}$ is the image embedding, and $E_{caption}$ is the caption embedding. The constant $c$ adjusts the weight of the caption embedding. This value is determined empirically through human evaluation of a small sample of generated questions. For the ShareGPT4V dataset (Chen et al., 2023), $c=0.2$ worked well, though this parameter may vary for different datasets depending on their quality and the descriptiveness of their captions to achieve optimal results.
This multimodal approach allows us to capture nuanced relationships between visual content and descriptive text, enhancing our ability to identify and group correlated images effectively. After constructing the multimodal embeddings, we applied UMAP (McInnes et al., 2018) to reduce the dimensionality of the vectors, projecting them into a lower-dimensional space for more efficient analysis and visualization, while preserving the essential structure of the data.
3.2 Grouping Images
We present two novel algorithms, Greedy Cluster Matching and Random Sampling with Iteration, designed to group correlated images prior to leveraging an open-source Large Language Model (LLM) for synthetic data generation in multi-image reasoning tasks. The emphasis on correlated images is crucial, as it facilitates challenging multi-image reasoning scenarios. These scenarios require the model to identify intricate relationships and differentiate between visually similar scenes, thus enhancing the complexity and realism of the reasoning process.
Figure 3: Left: Cluster Matching from Different Embedding Spaces, where images are sampled only from the union of clusters with the same color. Right: Sampling based on embedding distance for increased diversity from initial image (orange), where all images are considered within a single question.
Greedy Cluster Matching
After employing HDBSCAN (Malzer & Baum, 2020), a density-based clustering algorithm, to group the SigLIP and CLIP multimodal embeddings into coherent clusters, we developed a greedy algorithm to establish meaningful relationships between the two embedding spaces. Given cluster sets $C_{S}=\{S_{1},...,S_{m}\}$ from SigLIP and $C_{C}=\{C_{1},...,C_{n}\}$ from CLIP embeddings, the algorithm:
1) Orders clusters by size in descending order
2) Iteratively selects the largest remaining cluster $X_{max}=\arg\max_{X\in C_{S}\cup C_{C}}|X|$
3) For the selected cluster, finds its best match $Y_{best}$ from the other embedding space using: $Y_{best}=\arg\max_{Y}\frac{|X_{max}\cap Y|}{\frac{|X_{max}|+|Y|}{2}}$
This process continues until all clusters are matched or one set is exhausted, demonstrating high efficacy in producing quality image-caption pairs through dual-embedding validation. The score function, which measures overlap normalized by average cluster size, ensures each cluster corresponds to a semantically similar cluster from the other model. The detailed steps of this algorithm and pseudocode are presented in Figure 3 and Algorithm 1 (Appendix A.1).
While this approach effectively leverages both embedding models to confirm spatial relationships and associated semantic meanings, with images matched within clusters being corroborated by two independent embedding models, it has a notable limitation: sampling confined to matched clusters can lead to overly specialized image subjects, such as clustering only sheep-related images rather than diverse animal scenes, as demonstrated in Figure 6 (Appendix B.1). This specialization occurs because sampling is restricted to a single matched cluster, potentially limiting the diversity of selected images.
Random Sampling with Iteration
To address the limitations of cluster-confined sampling and transcend individual cluster boundaries, we developed Random Sampling with Iteration as our preferred solution. This method generates more generalized image subjects while maintaining semantic relationships through an iterative sampling approach. Given a set of embeddings $X=\{x_{1},...,x_{n}\}$, the algorithm:
1) Randomly selects an initial embedding index and initializes the selected set $S$
2) For each subsequent selection until reaching $N$ samples, computes a probability distribution over remaining embeddings based on: $p(x_{j})\propto\frac{1}{\sum_{u\in S}\|x_{j}-x_{u}\|^{k}+\epsilon}$, where $x_{j}\in X\setminus S$
3) Samples the next embedding according to this distribution, favoring points with smaller cumulative distances
Here, $k$ (default: 12) controls the emphasis on proximity - higher values more strongly favor nearby points by amplifying the differences in the distance calculations. The flexibility of this random vector sampling approach enables drawing from semantically related regions of the embedding space, while maintaining some randomness through probabilistic selection. This contributes to a more coherent dataset while still allowing for sufficient variety in the samples. For example, as shown in Figure 7 (Appendix B.2), this approach successfully samples related animal images while maintaining semantic coherence, in contrast to the disconnected clusters seen earlier. The detailed steps of this algorithm and pseudocode are presented in Figure 3 and Algorithm 2 (Appendix A.2).
By assembling related images before prompting the LLM, we create a more coherent and contextually rich input, enabling the model to generate more nuanced and relevant synthetic data for multi-image reasoning tasks.
3.3 Implementation
Our implementation of the SMiR dataset construction builds upon this approach of leveraging grouped images. After sampling correlated image-caption pairs, we developed a synthetic data generation process that incorporates these captions into system prompts for an open-source LLM. We specifically chose Meta Llama 3.1 70B Turbo for its significant efficiency advantages—up to 50 times cheaper and 10 times faster compared to GPT-4 (Kirkovska, 2024). This process generates complex multi-turn conversations between User and Assistant as seen in Table 2, with questions tailored to the selected images, as shown in Figure 6 and Figure 7 (Appendix B).
Our implementation employs two distinct types of system prompts to effectively capture different aspects of visual reasoning. The first type focuses on shorter visual questions, often involving OCR tasks, while the second addresses longer reasoning questions requiring in-depth analysis. To effectively capture the nuances of both complex reasoning scenarios and straightforward visual comprehension tasks, we developed separate, tailored prompts for each question type. These prompts were inspired by the methodologies of LLaVA (Liu et al., 2024b) and Chain-of-Thought (Wei et al., 2022), designed to generate multi-turn conversations that enhance the complexity and depth of interactions in our dataset. The detailed implementations of short-form prompts for visual tasks and long-form prompts for extended reasoning are presented in Appendices C.1 and C.2, respectively.
For the data source, we leveraged ShareGPT4V (Chen et al., 2023), an open-source dataset that combines data from LLaVA-Instruct (Liu et al., 2024a) and COCO (Lin et al., 2014). ShareGPT4V was initially developed using 100,000 captions generated by GPT4-Vision and subsequently expanded to 1.2 million image-caption pairs through an improved caption model, maintaining both quality and scale throughout the expansion process. To ensure diversity in our synthetic data, we processed the dataset in batches of 20,000 images, generating 5,000 synthetic conversations per batch. Each synthetic conversation contains an average of 4.65 images, and in total, we used 640,000 image-caption pairs to produce 160,000 synthetic conversations. We employed a vector sampling algorithm in conjunction with our long question prompts, with the primary objective of generating questions that were not only diverse but also challenged multi-step reasoning capabilities through multi-turn interactions. The modular design of our implementation allows for easy adaptation to other data resources in future applications.
Table 2: SMiR Dataset Statistics
Metric Value
Number of Samples 160,000
Maximum Number of Turns 24
Minimum Number of Turns 2
Average Number of Turns 9.65
Average Number of Images 4.65
Average User Tokens 25.51
Average Assistant Tokens 124.32
Open-Source LLM Meta Llama 3.1 70B Turbo
4 Multi-Image Benchmark
We create questions that challenge Vision Language Models (VLMs) to reason over multiple related images, analyze relationships, and derive meaning from series of images. The diverse range of tasks in this benchmark pushes the boundaries of multimodal understanding, requiring sophisticated analysis and interpretation across various visual contexts. Our benchmark compels models to provide detailed explanations for complex visual tasks, going beyond simple multiple-choice answers to demonstrate advanced reasoning capabilities. We implement an open-ended questionnaire with a judge model that evaluates responses in a pairwise structure, enabling relative comparisons between models and highlighting subtle differences in their ability to handle complex multi-image reasoning tasks.
4.1 Benchmark Overview
To address these challenges, we developed SMiR-Bench, a multi-turn benchmark consisting of 200 examples across seven diverse topics. These topics include:
  • Bird: identifying species and reasoning over distinguishing features
  • Matching: pairing photos based on visual cues
  • Object Character Recognition (OCR): reading and reasoning about academic texts
  • Pattern: identifying visual task patterns
  • Ranking: ordering everyday objects based on preferences
  • Storytelling: narrating based on image sequences, as shown in Figure 4
  • Visual: finding connections between images
The benchmark was curated collaboratively by a human annotator and GPT-4, utilizing images from the internet and the Shot2Story dataset (Han et al., 2023).
4.2 Evaluation Methodology
SMiR-Bench extends the Auto-Hard-Auto v0.1 (Li et al., 2024c) framework to the multimodal domain. The evaluation process begins with all models completing the benchmark tasks. Subsequently, one model is selected as the baseline, and a GPT-4o judge conducts pairwise comparisons between the baseline model’s responses and those of every other model. For each comparison, the judge evaluates responses on helpfulness, relevance, and conciseness. The judge’s assessments result in comparative rankings between two models (represented as A and B), such as $A\gg B$ (A significantly better than B), $A>B$ (A better than B), $A=B$ (A equal to B), or their inverses. Based on these pairwise comparisons and confidence levels, the framework generates an overall ranking of all models. This comprehensive approach enables a multi-turn, automatic, and challenging evaluation process, providing a nuanced assessment of model performance in multimodal tasks. By leveraging pairwise comparisons and confidence levels, it offers a clear view of how models perform relative to each other, complete with statistical reliability measures.
Figure 4: Evaluation Benchmark (Storytelling) Using GPT-4o as Judge
4.3 Results
To evaluate the effectiveness of our SMiR dataset, we finetune popular open-source VLMs and assess their performance on SMiR-Bench. Our experiments demonstrate significant improvements, with finetuned models achieving up to 8% better performance compared to their base versions.
4.4 Experiments
For our evaluation, we begin with two pre-trained models: Mantis-8B-siglip-llama3-pretrained (Jiang et al., 2024) and idefics-8b (Laurençon et al., 2024b). The first model combines a Llama-3-8B backbone with a SIGLIP encoder and is pretrained on CC3M (Gan et al., 2022). The second uses a Mistral-7B-v0.1 backbone with a perceiver (Jaegle et al., 2021) and MLP projection, pretrained on OBELICS. Both models were initially instruction-tuned on MANTIS-Instruct (721K examples) to create Mantis-8B-siglip-llama3 and Mantis-8B-Idefics2.
We then finetune these same pre-trained models on our SMiR dataset to create SMiR-8B-siglip-llama3-160 (trained on 160K examples) and SMiR-8B-Idefics2-100 (trained on 100K examples). To evaluate performance, we conduct pairwise comparisons using GPT-4 as a judge model against three specific baselines: GPT-4-Turbo to assess our models against a closed-source solution (Table 3), Mantis-8B-siglip-llama3 to evaluate SMiR-8B-siglip-llama3-160 (Table 4), and Mantis-8B-Idefics2 to compare with SMiR-8B-Idefics2-100 (Table 5). Remarkably, our SMiR-tuned models outperform their MANTIS-tuned counterparts despite being trained on a significantly smaller dataset. These results demonstrate the effectiveness of our SMiR dataset for finetuning, showing consistent performance improvements across all evaluation scenarios.
Table 3: SMiR-Bench Scores with GPT-4-Turbo Baseline
Model Name SMiR-Bench $\Delta$ 95% CI Average Tokens
Claude-3.5-Sonnet-20240620 54.2 (-4.0, 4.3) 362
GPT-4-Turbo 50.0 (0.0, 0.0) 359
Gemini-1.5-Pro 49.5 (-4.3, 4.5) 361
GPT-4o 46.3 (-4.0, 4.5) 316
Claude-3-Opus-20240229 34.4 (-4.6, 4.4) 321
SMiR-8B-siglip-llama3-160 5.0 +.8% (-2.2, 2.7) 183
Mantis-8B-siglip-llama3 4.2 (-1.2, 2.2) 146
SMiR-8B-Idefics2-100 4.5 +1.3% (-1.3, 1.4) 157
Mantis-8B-Idefics2 3.2 (-1.2, 1.2) 171
Idefics2-8B 2.5 (-1.0, 1.2) 118
Mantis-8B-siglip-llama3-pretrained 1.3 (-.8, .7) 207
LLaVA-v1.6-mistral-7b-hf 1.1 (-.7, .7) 317
Table 4: SMiR-Bench Scores with Mantis-8B-siglip-llama3 Baseline
Model Name SMiR-Bench $\Delta$ 95% CI Average Tokens
Claude-3-Opus-20240229 97.4 (-1.3, 1.0) 321
Claude-3-5-Sonnet-20240620 97.1 (-1.3, 1.2) 362
GPT-4-Turbo 96.4 (-1.8, 1.0) 359
Gemini-1.5-Pro 96.3 (-1.3, 1.3) 361
GPT-4o 91.5 (-2.4, 1.8) 316
SMiR-8B-siglip-llama3-160 58.1 +8.1% (-4.6, 5.3) 156
Mantis-8B-siglip-llama3 50.0 (0.0, 0.0) 146
LLaVA-v1.6-Mistral-7B-HF 15.8 (-2.4, 3.1) 317
Mantis-8B-siglip-llama3-pretraind 8.7 (-2.1, 2.1) 207
Table 5: SMiR-Bench Scores with Mantis-8B-Idefics2 Baseline
Model Name SMiR-Bench $\Delta$ 95% CI Average Tokens
Claude-3-Opus-20240229 98.1 (-1.2, 0.9) 321
Claude-3-5-Sonnet-20240620 98.1 (-1.1, 1.1) 362
Gemini-1.5-Pro 97.5 (-1.1, 0.9) 361
GPT-4-Turbo 95.9 (-1.5, 1.6) 359
GPT-4o 94.5 (-1.6, 1.5) 316
SMiR-8B-Idefics2-100 58.0 +8.0% (-5.3, 3.9) 157
Mantis-8B-Idefics2 50.0 (0.0, 0.0) 171
Idefics2-8B 29.7 (-3.7, 4.4) 118
LLaVA-v1.6-Mistral-7B-HF 16.7 (-3.4, 3.0) 317
Figure 5: SMiR Dataset Size vs. Benchmark Score
5 Conclusion
This paper introduces a synthetic data pipeline designed to enhance multi-image reasoning capabilities on open-source VLMs. By leveraging multimodal embeddings and grouping algorithms, the pipeline generates high-quality synthetic multi-image reasoning instruction tuning data. The approach yields up to 8% improvement on SMiR-Bench for popular open-source models, demonstrating the significant potential of synthetic data in advancing open-source VLM models.
Limitations
Our methods have several limitations. Random sampling with iteration is time-intensive due to the need for recalculating distance embeddings for each new image sampled. Further investigation is needed to determine the scalability of our synthetic data. Future research should focus on developing more time-efficient algorithms and optimizing data mixtures.
Broader Impact
This paper introduces a method for generating high-quality, cost-effective data for VLMs, addressing the growing challenge of data scarcity. By advancing these open-source techniques, we contribute to narrowing the performance gap between open and closed-source models, promoting more accessible and powerful multimodal AI.
Appendix A Algorithm Details
A.1 Greedy Cluster Matching Algorithm
We present the pseudocode for the Greedy Cluster Matching in Algorithm 1.
Let $C_{S}=\{S_{1},...,S_{m}\}$ and $C_{C}=\{C_{1},...,C_{n}\}$ be cluster sets from SigLIP and CLIP embeddings respectively. The algorithm proceeds as follows:
1. Select the largest cluster from either set: $X_{max}=\arg\max_{X\in C_{S}\cup C_{C}}|X|$
2. If $X_{max}\in C_{S}$, find the best match in $C_{C}$: $Y_{best}=\arg\max_{C_{j}\in C_{C}}score(X_{max},C_{j})$
3. If $X_{max}\in C_{C}$, find the best match in $C_{S}$: $Y_{best}=\arg\max_{S_{i}\in C_{S}}score(X_{max},S_{i})$
Where the score function is defined as:
$score(A,B)=\frac{|A\cap B|}{\frac{|A|+|B|}{2}}$
This process is repeated, greedily selecting the largest remaining cluster and finding its best match, until all clusters are matched or one set is exhausted.
Algorithm 1 Greedy Cluster Matching Algorithm
1:Two lists of clusters $c1$ and $c2$
2:List of matched cluster pairs
[…]
7:while $c1$ is not empty and $c2$ is not empty do
8:     if $\text{len}(c1[0])\geq\text{len}(c2[0])$ then
[…]
11:     else
[…]
14:     end if
[…]
24:         end if
25:     end for
26:     if best_match is not None then
[…]
32:     end if
33:end while
[…]
A.2 Random Sampling with Iteration
We present the pseudocode for the Random Sampling with Iteration in Algorithm 2.
Let $X=\{x_{1},...,x_{n}\}$ be the set of embeddings.
$k$ is a parameter that determines the power of the distance calculation (default to 12), and $N$ is the desired number of selected embeddings.
1. Randomly select an initial embedding: $s_{1}\in X$
2. Initialize selected set $S=\{s_{1}\}$
3. For $i=2$ to $N$: $s_{i}=\arg\max_{x_{j}\in X\setminus S}\sum_{u\in S}\|x_{j}-x_{u}\|^{k}$, $S=S\cup\{s_{i}\}$
4. Return $S$
This formulation captures the process of iteratively selecting embeddings based on their cumulative distance from all previously selected embeddings, raised to the power $k$.
Algorithm 2 Random Sampling with Iteration
1:
2:$X$: Set of embeddings
3:$N$: Number of samples to select
4:$k$: Power factor for distance calculation (default: 12)
5:Set of selected indices
6:$selected\leftarrow[]$
7:$n\leftarrow|X|$triangleright Number of embeddings
8:for $i=1$ to $N$ do
9:     $distances\leftarrow\text{zeros}(n)$
10:     if $selected$ is empty then
11:         $sampled\_index\leftarrow\text{random\_integer}(0,n-1)$
12:     else
13:         for $j=0$ to $n-1$ do
14:              if $j\in selected$ then
15:                  $distances[j]\leftarrow\infty$
16:              else
17:                  $distances[j]\leftarrow\sum_{u\in selected}\|\|X[j]-X[u]\|\|^{k}$
18:              end if
19:         end for
20:         $inverted\_distances\leftarrow\frac{1}{distances+\epsilon}$triangleright $\epsilon$ is a small constant
21:         $distribution\leftarrow\frac{inverted\_distances}{\sum inverted\_distances}$
22:         $sampled\_index\leftarrow\text{random\_choice}(\text{range}(n),p=distribution)$
23:     end if
24:     $selected.\text{append}(sampled\_index)$
25:end for
26:return $selected$
Appendix B Data Samples
For the sake of brevity, we have included only two examples from the multi-image data we have generated, featuring questions with a related theme from both the Greedy Cluster Matching Algorithm and Random Sampling with Iteration, highlighting how the algorithms differ in their approach to generating questions.
B.1 Greedy Cluster Matching Algorithm
Samples obtained through Greedy Cluster Matching typically feature similar subjects and shot compositions, but when paired with carefully crafted prompts, these similarities can be leveraged to generate more challenging and nuanced questions.
Figure 6: Images sampled from the same matched cluster often feature similar subjects or scenes.
B.2 Random Sampling with Iteration
Random sampling tends to yield greater diversity in subjects compared to Greedy Clustering, striking a balance between variety and relatedness that can potentially lead to more robust and wide-ranging question sets.
Figure 7: Images sampled using the iterative algorithm allow for different yet related subjects (e.g., various animal species)
Appendix C Prompt
While prompts play a crucial role in data generation, optimizing them remains a significant challenge. After numerous iterations, we identified two particularly effective prompts for multi-image data generation.
C.1 LLaVA Style Prompt
Inspired by LLaVA (Liu et al., 2024b), our approach utilizes a specialized prompt to address simpler multi-image and single-image tasks, focusing on more straightforward visual comprehension and analysis.
You are an AI visual assistant that can analyze multiple images. You receive four to five images with their corresponding captions in an array. The task is to use the provided images and captions to create a challenging and complex question that involves comparison, ranking, storytelling, or logical reasoning across the images and then provide a detailed answer. The question should require a deep understanding of the visual content and advanced reasoning based on that content.
Create questions that ask to compare elements across the images, such as identifying which image best represents a critical or turning point moment, quality, or characteristic; formulate questions that require ranking the images based on intricate and plausible criteria, such as strategic importance, sequence, or visual impact; develop questions that involve piecing together a narrative from the images, understanding a sophisticated sequence of events, or explaining a complex progression shown; and ask questions that require advanced logical reasoning to deduce why certain elements are present, the purpose behind actions shown, or the broader implications of what is depicted.
Frame a question that requires advanced analyzing and reasoning about the images and their captions, and provide a detailed answer based on the visual content and captions, explaining the reasoning process and the conclusions drawn. Also, provide 3-4 follow-up questions to deepen the analysis based on the potential responses, and provide detailed answers for the follow-up questions as well.
Example questions include: - ‘‘Which image best represents the pivotal turning point of the event, and why?’’ - ‘‘Rank the images based on the strategic importance of the actions shown, from highest to lowest.’’ - ‘‘How do the images collectively tell the intricate story of the event, and what can be inferred about the key strategic moments?’’ - ‘‘What could be the underlying reasons behind the specific actions taken in each image, and how do they relate to the overall context in a broader sense?’’
[…]
Return your challenging and complex question and follow-up questions in the format: ‘‘User: [INSERT QUESTION], Assistant: [INSERT ANSWER] User: [INSERT FOLLOW-UP QUESTION 1], Assistant: [INSERT FOLLOW-UP ANSWER 1] User: [INSERT FOLLOW-UP QUESTION 2], Assistant: [INSERT FOLLOW-UP ANSWER 2] User: [INSERT FOLLOW-UP QUESTION 3], Assistant: [INSERT FOLLOW-UP ANSWER 3]...’’
C.2 Longer Prompt
Our approach aims to generate more complex, multi-turn questions that require in-depth reasoning across multiple images.
You are an AI visual assistant capable of analyzing multiple images, including both visual content and textual elements using Optical Character Recognition (OCR). You will receive four to five images, each potentially accompanied by captions and containing text, numbers, signs, or other recognizable characters. Your task is to create a plausible and challenging question that involves comparison, ranking, storytelling, logical reasoning, or detailed textual analysis across the images, and then provide a detailed answer.
For Visual Analysis: Frame questions that require understanding and reasoning about the visual content, such as comparing elements across images, identifying which image best represents a specific moment, quality, or characteristic, ranking the images based on logical and plausible criteria (e.g., importance, sequence, or visual quality), or piecing together a narrative from the images to explain a sequence of events or the progression shown. Your questions should encourage deep engagement with the visual content and require advanced reasoning or interpretation.
For OCR-Based Analysis: Frame questions that require detailed analysis of the textual content in the images, such as counting specific items or words, identifying differences or similarities between the images, verifying the accuracy of information, or deducing logical conclusions based on the textual data. Your questions should encourage users to engage deeply with the textual content and require precise reasoning or interpretation.
Question Creation: Create a question that requires analyzing and reasoning about the images and their captions or textual elements. Provide a detailed answer based on the visual content, captions, and/or OCR analysis, explaining the reasoning process and conclusions drawn.
Follow-Up Questions: Additionally, include 3-4 follow-up questions that delve deeper into the analysis based on the initial question. Provide detailed answers for each follow-up question, further expanding on the reasoning and conclusions.
Example Questions: - ‘‘Which image best represents the climax of the event, and why?’’ - ‘‘Rank the images based on the level of engagement of the individuals shown, from highest to lowest.’’ - ‘‘How many times does the word \‘‘urgent\’’ appear across all images, and in which image is it most prominently displayed?’’ - ‘‘Identify whether all the images contain the same warning label. If there are differences, describe them.’’
Example Follow-Up Questions: - ‘‘Why do you think the climax identified in Image (image number) was critical to the outcome of the event?’’ - ‘‘How do the textual differences in Image (image number) affect your understanding of the event’s context?’’ - ‘‘What additional visual details in Image (image number) support your interpretation of the story?’’ - ‘‘How does the sequence of numbers in these images relate to the broader narrative depicted?’’
Format: Return your challenging and complex question, detailed answer, and follow-up questions in the following format: ‘‘User: [INSERT QUESTION], Assistant: [INSERT ANSWER] User: [INSERT FOLLOW-UP QUESTION 1], Assistant: [INSERT FOLLOW-UP ANSWER 1] User: [INSERT FOLLOW-UP QUESTION 2], Assistant: [INSERT FOLLOW-UP ANSWER 2] User: [INSERT FOLLOW-UP QUESTION 3], Assistant: [INSERT FOLLOW-UP ANSWER 3]...’’

### v2

@@
 Abstract
-Vision-Language Models (VLMs) have shown strong performance in understanding single images, aided by numerous high-quality instruction datasets.
-However, multi-image reasoning tasks are still under-explored in the open-source community due to two main challenges: (1) scaling datasets with multiple correlated images and complex reasoning instructions is resource-intensive and maintaining quality is difficult, and (2) there is a lack of robust evaluation benchmarks for multi-image tasks.
-To address these issues, we introduce SMiR, an efficient synthetic data-generation pipeline for multi-image reasoning, and a high-quality dataset generated using this pipeline.
-Our pipeline efficiently extracts highly correlated images using multimodal embeddings, combining visual and descriptive information and leverages open-source LLMs to generate quality instructions.
-Using this pipeline, we generated 160K synthetic training samples, offering a cost-effective alternative to expensive closed-source solutions.
-Additionally, we present SMiR-Bench, a novel multi-image reasoning evaluation benchmark comprising 200 diverse examples across 7 complex multi-image reasoning tasks.
-SMiR-Bench is multi-turn and utilizes a VLM judge to evaluate free-form responses, providing a comprehensive assessment of model expressiveness and reasoning capability across modalities.
-We demonstrate the effectiveness of SMiR dataset by fine-tuning several open-source VLMs and evaluating their performance on SMiR-Bench.
-Our results show that models trained on our dataset outperform baseline models in multi-image reasoning tasks up to 8% with a much more scalable data pipeline.
+Vision-Language Models (VLMs) excel at understanding single images, aided by high-quality instruction datasets.
+However, multi-image reasoning remains underexplored in the open-source community due to two key challenges: (1) scaling datasets with correlated images and complex reasoning instructions is resource-intensive, and (2) robust evaluation benchmarks for multi-image tasks are lacking.
+To address this, we introduce SMiR, a synthetic data-generation pipeline for multi-image reasoning, along with a high-quality dataset generated using this pipeline.
+SMiR efficiently extracts correlated images via multimodal embeddings, integrates visual and descriptive information, and leverages open-source LLMs to generate quality instructions.
+Using this approach, we produce 160K synthetic training samples, offering a cost-effective alternative to closed-source solutions.
+Additionally, we present SMiR-Bench, a multi-image reasoning benchmark comprising 200 diverse examples across seven complex reasoning tasks.
+SMiR-Bench is multi-turn and employs a VLM judge to evaluate free-form responses, providing a comprehensive assessment of model expressiveness and reasoning capability across modalities.
+We demonstrate the effectiveness of SMiR by fine-tuning open-source VLMs and evaluating them on SMiR-Bench.
 Our codebase is available at https://github.com/togethercomputer/SMiR.
 1 Introduction
-Vision-Language Models (VLMs) have shown impressive capabilities in tasks involving single images, particularly open-source models that have benefited from high-quality instruction datasets (Laurençon et al., 2024b; Zhang et al., 2023; Xu et al., 2022).
-However, in multi-image tasks, such as comparing or analyzing relationships between multiple images, the performance of open-source VLMs (Liu et al., 2024b; a; Li et al., 2024a; Awadalla et al., 2023; Yao et al., 2024; Wang et al., 2023b) significantly lags behind that of their closed-source counterparts like GPT-4 (Achiam et al., 2023), Claude 3.5 Sonnet (Anthropic, 2024a), Claude 3 (Anthropic, 2024b), and Gemini 1.5 (Reid et al., 2024).
-While we do not have full visibility into how open-source models are trained, a reasonable assumption for their underperformance in multi-image reasoning tasks is the advantage closed-source models have in terms of training datasets.
-Closed-source developers likely benefit from access to large, high-quality datasets, which are critical for training advanced models capable of reasoning across multiple images.
-In contrast, open-source models may face limitations in accessing or constructing datasets that are specifically tailored to these tasks, which impacts their overall performance.
-One major challenge lies in collecting and curating large-scale datasets with highly correlated images.
-Identifying meaningful connections between images requires vast amounts of data and sophisticated algorithms, which most multi-image instruction-tuning datasets lack.
-For example, MANTIS (Jiang et al., 2024) often pairs unrelated images in the same reasoning task, unintentionally simplifying the challenge.
-Consider a question like: Compare the behaviors and characteristics shown in these images.
-How do they demonstrate different aspects of survival and adaptation?
-Previous methods may match it with a relevant image of a dog alongside completely unrelated images like a drone hovering in the air and shipping containers at a port.
-While the question naturally applies to analyzing an animal’s behavioral adaptations, its pairing with technological and industrial objects (drones, logistics, etc.) forces superficial comparisons that dilute the biological focus of the inquiry.
-This makes the task easier, as the unrelated items can only be discussed through loose metaphorical connections to survival and adaptation.
-Figure 1: Comparison of top (example of MANTIS dataset sharegpt4v-pt_125, which can have a data sample with animals, but are concatenated together with unrelated images) vs.
-bottom (SMiR dataset where related images are together) for multi-image reasoning.
-In contrast, comparing related images—like a series of different animals in their natural habitats—demands a deeper understanding of predator behavior and ecological adaptation.
-Here, the task requires a more nuanced comparison of hunting techniques, territorial behaviors, and social structures within the same family of species.
-MMDU-45K (Liu et al., 2024e) attempts to address this by clustering image captions and category tags, but it overlooks the crucial visual relationships.
-These shortcomings highlight the need for datasets with truly correlated images to foster complex multi-image reasoning tasks that better reflect real-world analytical challenges.
-Second, scaling up the number of highly correlated images presents significant challenges.
-Existing datasets such as MANTIS, MMDU-45K, and MultiInstruct (Xu et al., 2022) require extensive human curation and annotation, resulting in a labor-intensive and time-consuming process.
-To minimize human effort, researchers have leveraged GPT-4 family models (Peng et al., 2023; Wang et al., 2023a) to generate synthetic datasets—including MMInstruct (Liu et al., 2024d), Multimodal ArXiv (Li et al., 2024b), MIMIC-IT (Li et al., 2023a), StableLLaVA-Instruct (Li et al., 2023c), and SVIT-Instruct (Zhao et al., 2023).
-However, this method proves expensive and difficult to scale effectively.
-Third, evaluating multi-image reasoning is challenging due to the limitations of traditional assessment methods.
-Multiple-choice formats, used in previous multi-image reasoning benchmarks (Meng et al., 2024; Fu et al., 2024; Wang et al., 2024a; Yue et al., 2024), fail to capture the nuanced reasoning process and justifications behind a model’s answers.
-This is problematic because models can achieve similar performance on multiple-choice tasks while having vastly different reasoning capabilities.
-In contrast, free-form responses allow models to demonstrate their thought processes, providing a more comprehensive evaluation of their true reasoning abilities and distinguishing between models that rely on superficial pattern matching and those with a deeper understanding.
-To address these challenges, we propose a synthetic data generation pipeline, SMiR, for multi-image reasoning, along with a human-annotated evaluation benchmark, SMiR-Bench.
-SMiR generates correlated and challenging multi-image reasoning questions, while SMiR-Bench evaluates models using free-response tasks in complex multi-image scenarios.
-Our contributions are as follows:
-  • We present SMiR, a pipeline for multi-image instruction tuning.
-To construct SMiR, we apply existing algorithms in a novel way to collect multiple images with high correlation.
-Our approach combines cluster sampling for data quality robustness and graph iteration sampling for diversity.
-Both algorithms leverage multimodal embeddings that integrate image and caption information to effectively group correlated images, enabling challenging multi-image instruction tuning scenarios.
-  • We develop a scalable synthetic multimodal data generation pipeline that leverages open-source LLMs such as Meta Llama 3.1 70B Instruct Turbo (Dubey et al., 2024).
-This approach eliminates the need for expensive closed-source models, reducing costs by up to 50 times and increasing speed by up to 10 times (Kirkovska, 2024), while significantly minimizing human annotation efforts.
-  • We present a new multimodal evaluation benchmark that uses multi-turn interactions and free-form responses to assess both final answers and reasoning processes in complex multi-image tasks.
-The evaluation is conducted pairwise, directly comparing model performances.
-Using reference answers from GPT-4-Turbo and other open-source models as baselines, we observe up to an 8% improvement in win rate for instruction-tuned models trained on the SMiR dataset.
-2 Related Works
-Vision Language Models
-We focus on instruction tuning Vision-Language Models (VLMs) that utilize a pretrained Large Language Model (LLM) backbone because this approach is cost-effective and more accessible for the open-source community.
-Since the backbone responsible for language understanding is already trained, the overall training process becomes simpler and requires fewer resources.
-Our primary task involves aligning the vision encoder—typically architectures like Vision Transformer (ViT) (Dosovitskiy, 2020), SigLIP (Zhai et al., 2023), or CLIP (Radford et al., 2021)—with the LLM backbone.
-This alignment is facilitated through linear layers that connect the vision encoder to the backbone, enabling the integration of visual and textual information.
-For instance, BLIP-2 (Li et al., 2023b) uses OPT (Zhang et al., 2022) and FLAN-T5 (Chung et al., 2022) as backbones, MiniGPT-4 (Zhu et al., 2023) utilizes Vicuna (Chiang et al., 2023), and Qwen-2-VL (Wang et al., 2024b) employs Qwen-2-1.5B (Yang et al., 2024) as the language backbone.
-In this paper, we focus on creating a high-quality multi-image reasoning dataset for instruction tuning instead of large-scale interleaved pretraining datasets like OBELICS (Laurençon et al., 2024a), MINT-1T (Awadalla et al., 2024), and LAION-5B (Schuhmann et al., 2022).
-Multi-Image Reasoning Data
-Recent advancements in multi-image reasoning instruction tuning datasets include MANTIS (Jiang et al., 2024) and MMDU-45K (Liu et al., 2024e), both aiming to improve reasoning capabilities in VLMs.
-However, these datasets have limitations in their approaches.
-MANTIS randomly concatenates single image pairs from LLaVA-665k (Liu et al., 2024a), which often results in uncorrelated images within multi-image scenarios, potentially undermining the complexity of reasoning tasks.
-MMDU-45K attempts to address this issue by utilizing sentence transformers (Reimers, 2019) with description text and clustering techniques to group related images, but does not consider visual components.
-The dataset is then further enhanced, assisted by GPT-4 to generate comprehensive answers for the grouped images.
-Building upon these efforts, SMiR introduces a novel approach that leverages both vital visual and caption information to ensure highly correlated images within multi-image sets with the use of open-source LLMs.
-These scalable methods leads to the generation of more challenging questions that require deeper analysis and understanding of visual relationships, pushing the boundaries of multi-image reasoning capabilities in VLMs.
+Vision-Language Models (VLMs) have demonstrated strong performance on single-image tasks, particularly open-source models trained with high-quality instruction datasets (Laurençon et al., 2024b; Zhang et al., 2023; Xu et al., 2022).
+However, open-source VLMs struggle with multi-image reasoning tasks—such as comparing and analyzing relationships between multiple images—lagging behind closed-source models like GPT-4 (Achiam et al., 2023), Claude 3.5 Sonnet (Anthropic, 2024), and Gemini 1.5 (Reid et al., 2024).
+This performance gap likely stems from the superior multi-image datasets available to closed-source models, which are essential for effective reasoning across multiple images.
+A key challenge in multi-image reasoning is collecting and curating large-scale datasets with highly correlated images.
+Existing datasets such as MANTIS (Jiang et al., 2024) often pair unrelated images, simplifying the task and limiting the complexity of reasoning.
+MMDU-45K (Liu et al., 2024f) improves upon this by clustering images based on captions and category tags but overlooks purely visual relationships.
+Consequently, datasets that explicitly ensure image correlation are needed to advance complex multi-image reasoning.
+Scaling such datasets further exacerbates the challenge.
+Human curation, as required in datasets like MultiInstruct (Xu et al., 2022), is resource-intensive.
+While synthetic datasets generated via GPT-4 (Peng et al., 2023; Wang et al., 2023) offer a scalable alternative (Liu et al., 2024e; Li et al., 2024a; 2023a), they remain expensive and difficult to scale effectively.
+Another challenge is evaluation.
+Existing multi-image benchmarks, such as MMIU (Meng et al., 2024) and MMMU (Yue et al., 2024), rely on multiple-choice questions, failing to assess the depth of a model’s reasoning.
+While some benchmarks incorporate free-form responses (Liu et al., 2024c), they lack multi-turn interactions, which are essential for evaluating complex reasoning processes.
+To address these issues, we introduce SMiR, a synthetic data generation pipeline for multi-image reasoning, and SMiR-Bench, a multi-image evaluation benchmark. Our key contributions are:
+  • We propose SMiR, a pipeline that efficiently identifies correlated images using multimodal embeddings, applying cluster sampling and graph iteration sampling to ensure diversity and robustness in instruction tuning.
+  • We develop a scalable synthetic data generation pipeline using open-source LLMs such as Meta Llama 3.1 70B Instruct Turbo (Dubey et al., 2024), reducing costs by up to 50x and increasing speed by up to 10x compared to closed-source alternatives (Kirkovska, 2024).
+  • We introduce SMiR-Bench, a multi-turn evaluation benchmark that assesses models via free-form responses and reasoning justifications, using GPT-4-Turbo and other baselines for comparison.
+Models trained on SMiR improve win rates by up to 8% over existing instruction-tuned models.
+By advancing both dataset generation and evaluation, SMiR facilitates the development of stronger open-source VLMs for multi-image reasoning.
+Table 1: Comparison of datasets highlighting key characteristics and methodologies.
 Datasets Multimodal Embedding Correlated Images No Human-Annotation Open-Source LLM
-Mantis
-MMDU
+Mantis Jiang et al. (2024)
+MMDU Liu et al. (2024f)
 SMiR
-Table 1: Comparison of datasets highlighting key characteristics and methodologies.
-Multi-Image Reasoning Benchmarks
-Recent VLM benchmarks (Chiang et al., 2024; Lin et al., 2024; Liu et al., 2024c) have made strides by incorporating free-response evaluations, marking a significant improvement over traditional multiple-choice formats.
-However, these benchmarks still lack a comprehensive approach that combines automatic, multi-turn, and pairwise evaluation capabilities.
-Our benchmark addresses this gap, drawing inspiration from Auto-Hard-Auto v0.1 (Li et al., 2024c).
-We have adapted and expanded this framework to enable robust multimodal evaluation, providing a more holistic assessment of VLM performance across complex, multi-image reasoning tasks.
-This approach allows for a deeper analysis of both the final answers and the underlying reasoning processes employed by VLMs in real-world SMiR-Bench scenarios.
-Figure 2: Our end-to-end pipeline: from sampling and LLM prompting to conversation generation.
-The example conversation is based on a sports scenario, demonstrating the pipeline’s ability to generate contextually relevant multi-turn dialogues.
+Figure 1: Our end-to-end pipeline converts image-caption pairs into synthetic multi-turn conversations using multimodal embeddings, strategic sampling, and LLM prompting.
+The example, based on a sports scenario, illustrates how the pipeline generates contextually rich dialogues by leveraging visual relationships.
+Figure 2: Comparison of an example from the MANTIS dataset, which includes samples from ShareGPT4V-PT where unrelated images such as animals and other objects are concatenated (top), vs.
+an example from the SMiR dataset, where only related images are grouped together (bottom), for multi-image reasoning.
 2 SMiR: Synthetic Multi-Image Reasoning Data Pipeline
-SMiR pipeline efficiently generates complex multi-image reasoning data through a carefully designed methodology that reflects key design considerations and trade-offs.
 Starting with a large-scale image-caption dataset $D$ containing $N$ image-caption pairs $(I_{i},C_{i})_{i=1}^{N}$, SMiR first constructs multimodal embeddings for each pair.
-Then, grouping algorithms identify correlations between these embeddings.
+Then, sampling algorithms identify correlations between these embeddings.
 Finally, open-source LLMs are used to generate complex question-answer pairs based on the sampled correlations.
-In the following subsections, we describe our approach’s core components—the construction of multimodal embeddings, grouping algorithms, and synthetic data generation implementation—while examining the critical design decisions behind our sampling algorithms, prompts, and data sources that shaped the final SMiR dataset.
+The end-to-end pipeline is shown in Fig.
+1.
 2.1 Multimodal Embedding Construction
-While embedding-based approaches have been used before to capture relationships between images and text, relying solely on either image or caption embeddings limits our ability to fully understand the connections between visual content and descriptive text.
-A more comprehensive approach requires incorporating both contexts simultaneously.
-To address this, we developed a method that combines image and caption information into a single multimodal embedding.
-Our approach utilizes SigLIP or CLIP embeddings for the images, combined with corresponding caption embeddings.
-We formulated the multimodal embedding as follows:
-$E_{multimodal}=E_{image}+c\cdot E_{caption}$ (1)
-where $E_{multimodal}$ is the multimodal embedding, $E_{image}$ is the image embedding, and $E_{caption}$ is the caption embedding.
-The constant $c$ adjusts the weight of the caption embedding.
-This value is determined empirically through human evaluation of a small sample of generated questions.
-For the ShareGPT4V dataset (Chen et al., 2023), $c=0.2$ worked well, though this parameter may vary for different datasets depending on their quality and the descriptiveness of their captions to achieve optimal results.
-This multimodal approach allows us to capture nuanced relationships between visual content and descriptive text, enhancing our ability to identify and group correlated images effectively.
-After constructing the multimodal embeddings, we applied UMAP (McInnes et al., 2018) to reduce the dimensionality of the vectors, projecting them into a lower-dimensional space for more efficient analysis and visualization, while preserving the essential structure of the data.
-3.2 Grouping Images
-We present two novel algorithms, Greedy Cluster Matching and Random Sampling with Iteration, designed to group correlated images prior to leveraging an open-source Large Language Model (LLM) for synthetic data generation in multi-image reasoning tasks.
-The emphasis on correlated images is crucial, as it facilitates challenging multi-image reasoning scenarios.
-These scenarios require the model to identify intricate relationships and differentiate between visually similar scenes, thus enhancing the complexity and realism of the reasoning process.
-Figure 3: Left: Cluster Matching from Different Embedding Spaces, where images are sampled only from the union of clusters with the same color.
-Right: Sampling based on embedding distance for increased diversity from initial image (orange), where all images are considered within a single question.
-Greedy Cluster Matching
-After employing HDBSCAN (Malzer & Baum, 2020), a density-based clustering algorithm, to group the SigLIP and CLIP multimodal embeddings into coherent clusters, we developed a greedy algorithm to establish meaningful relationships between the two embedding spaces.
-Given cluster sets $C_{S}=\{S_{1},...,S_{m}\}$ from SigLIP and $C_{C}=\{C_{1},...,C_{n}\}$ from CLIP embeddings, the algorithm:
-1) Orders clusters by size in descending order
-2) Iteratively selects the largest remaining cluster $X_{max}=\arg\max_{X\in C_{S}\cup C_{C}}|X|$
-3) For the selected cluster, finds its best match $Y_{best}$ from the other embedding space using: $Y_{best}=\arg\max_{Y}\frac{|X_{max}\cap Y|}{\frac{|X_{max}|+|Y|}{2}}$
-This process continues until all clusters are matched or one set is exhausted, demonstrating high efficacy in producing quality image-caption pairs through dual-embedding validation.
-The score function, which measures overlap normalized by average cluster size, ensures each cluster corresponds to a semantically similar cluster from the other model.
-The detailed steps of this algorithm and pseudocode are presented in Figure 3 and Algorithm 1 (Appendix A.1).
-While this approach effectively leverages both embedding models to confirm spatial relationships and associated semantic meanings, with images matched within clusters being corroborated by two independent embedding models, it has a notable limitation: sampling confined to matched clusters can lead to overly specialized image subjects, such as clustering only sheep-related images rather than diverse animal scenes, as demonstrated in Figure 6 (Appendix B.1).
-This specialization occurs because sampling is restricted to a single matched cluster, potentially limiting the diversity of selected images.
-Random Sampling with Iteration
-To address the limitations of cluster-confined sampling and transcend individual cluster boundaries, we developed Random Sampling with Iteration as our preferred solution.
-This method generates more generalized image subjects while maintaining semantic relationships through an iterative sampling approach.
-Given a set of embeddings $X=\{x_{1},...,x_{n}\}$, the algorithm:
-1) Randomly selects an initial embedding index and initializes the selected set $S$
-2) For each subsequent selection until reaching $N$ samples, computes a probability distribution over remaining embeddings based on: $p(x_{j})\propto\frac{1}{\sum_{u\in S}\|x_{j}-x_{u}\|^{k}+\epsilon}$, where $x_{j}\in X\setminus S$
-3) Samples the next embedding according to this distribution, favoring points with smaller cumulative distances
-Here, $k$ (default: 12) controls the emphasis on proximity - higher values more strongly favor nearby points by amplifying the differences in the distance calculations.
-The flexibility of this random vector sampling approach enables drawing from semantically related regions of the embedding space, while maintaining some randomness through probabilistic selection.
-This contributes to a more coherent dataset while still allowing for sufficient variety in the samples.
-For example, as shown in Figure 7 (Appendix B.2), this approach successfully samples related animal images while maintaining semantic coherence, in contrast to the disconnected clusters seen earlier.
-The detailed steps of this algorithm and pseudocode are presented in Figure 3 and Algorithm 2 (Appendix A.2).
-By assembling related images before prompting the LLM, we create a more coherent and contextually rich input, enabling the model to generate more nuanced and relevant synthetic data for multi-image reasoning tasks.
-3.3 Implementation
-Our implementation of the SMiR dataset construction builds upon this approach of leveraging grouped images.
-After sampling correlated image-caption pairs, we developed a synthetic data generation process that incorporates these captions into system prompts for an open-source LLM.
-We specifically chose Meta Llama 3.1 70B Turbo for its significant efficiency advantages—up to 50 times cheaper and 10 times faster compared to GPT-4 (Kirkovska, 2024).
-This process generates complex multi-turn conversations between User and Assistant as seen in Table 2, with questions tailored to the selected images, as shown in Figure 6 and Figure 7 (Appendix B).
-Our implementation employs two distinct types of system prompts to effectively capture different aspects of visual reasoning.
-The first type focuses on shorter visual questions, often involving OCR tasks, while the second addresses longer reasoning questions requiring in-depth analysis.
-To effectively capture the nuances of both complex reasoning scenarios and straightforward visual comprehension tasks, we developed separate, tailored prompts for each question type.
-These prompts were inspired by the methodologies of LLaVA (Liu et al., 2024b) and Chain-of-Thought (Wei et al., 2022), designed to generate multi-turn conversations that enhance the complexity and depth of interactions in our dataset.
-The detailed implementations of short-form prompts for visual tasks and long-form prompts for extended reasoning are presented in Appendices C.1 and C.2, respectively.
-For the data source, we leveraged ShareGPT4V (Chen et al., 2023), an open-source dataset that combines data from LLaVA-Instruct (Liu et al., 2024a) and COCO (Lin et al., 2014).
-ShareGPT4V was initially developed using 100,000 captions generated by GPT4-Vision and subsequently expanded to 1.2 million image-caption pairs through an improved caption model, maintaining both quality and scale throughout the expansion process.
-To ensure diversity in our synthetic data, we processed the dataset in batches of 20,000 images, generating 5,000 synthetic conversations per batch.
-Each synthetic conversation contains an average of 4.65 images, and in total, we used 640,000 image-caption pairs to produce 160,000 synthetic conversations.
-We employed a vector sampling algorithm in conjunction with our long question prompts, with the primary objective of generating questions that were not only diverse but also challenged multi-step reasoning capabilities through multi-turn interactions.
-The modular design of our implementation allows for easy adaptation to other data resources in future applications.
-Table 2: SMiR Dataset Statistics
+Embedding-based approaches have been widely used to model relationships between images and text.
+However, relying solely on image or caption embeddings limits the ability to fully capture the connections between visual content and descriptive text.
+A more comprehensive approach requires integrating both modalities.
+To achieve this, we construct a multimodal embedding by combining SigLIP or CLIP image embeddings with their corresponding caption embeddings:
+$E_{\text{multimodal}}=E_{\text{image}}+c\cdot E_{\text{caption}}$ (1)
+where $E_{\text{multimodal}}$ represents the fused embedding, and $c$ adjusts the weight of the caption embedding.
+This parameter is tuned empirically based on small-scale human evaluations.
+For the ShareGPT4V dataset (Chen et al., 2023), $c=0.2$ provided the best balance between visual and textual contributions.
+The optimal value may vary depending on dataset quality and caption descriptiveness.
+This multimodal approach enables better identification and clustering of correlated images by capturing both visual and textual relationships.
+We apply UMAP (McInnes et al., 2018) for dimensionality reduction, preserving essential data structures while facilitating efficient analysis and visualization.
+2.2 Random Sampling with Iteration
+After constructing multimodal embeddings, we apply a random sampling algorithm to select semantically related images, ensuring diversity while preserving coherence.
+Given a set of embeddings $X=\{x_{1},...,x_{n}\}$, the algorithm proceeds as follows:
+  1. 1.
+    Randomly selects an initial embedding and initializes the selected set $S$.
+  2. 2.
+    Iteratively selects embeddings until reaching $N$ samples, using the probability distribution:
+    $p(x_{j})\propto\frac{1}{\sum_{u\in S}\|x_{j}-x_{u}\|^{k}+\epsilon},\quad x_{j}\in X\setminus S$ (2)
+  3. 3.
+    Samples the next embedding based on this probability, favoring closer points.
+Here, $k$ (default: 12) controls the emphasis on proximity—higher values prioritize nearby points more strongly by amplifying distance differences.
+This approach balances semantic coherence and diversity, allowing for meaningful multi-image relationships while introducing some randomness.
+For example, as shown in Fig.
+6 (Section C.2), this method effectively clusters related animal images while avoiding the disconnected groupings seen in simpler sampling methods.
+The full algorithm and pseudocode are provided in Algorithm 2 (Section B.2).
+By assembling related images before prompting the LLM, we generate richer contextual inputs, enabling the model to produce more nuanced and relevant synthetic data for multi-image reasoning tasks.
+Table 2: Statistics of the SMiR dataset
 Metric Value
@@
 Open-Source LLM Meta Llama 3.1 70B Turbo
+2.3 Implementation
+Our implementation of the SMiR dataset construction builds on the structured sampling of correlated image-caption pairs.
+After grouping related images, we generate synthetic data by incorporating these captions into system prompts for an open-source LLM.
+We use Meta Llama 3.1 70B Turbo due to its efficiency, as it is up to 50 times cheaper and 10 times faster than GPT-4 (Kirkovska, 2024).
+This process produces complex, multi-turn conversations between the user and assistant, as shown in Table 2.
+Sampled images and corresponding generated questions are illustrated in Fig.
+5 and Fig.
+6 (Appendix C).
+To capture different aspects of visual reasoning, we design two types of system prompts.
+The first type focuses on shorter visual questions, often requiring OCR-based understanding.
+The second type is designed for long-form reasoning, where multi-step inference and deeper contextual understanding are needed.
+These prompts are inspired by LLaVA (Liu et al., 2024b) and Chain-of-Thought reasoning (Wei et al., 2022), enabling the generation of multi-turn conversations that enhance dataset complexity.
+The detailed structures of these prompts are provided in Section D.1 and Section D.2.
+For data sources, we use ShareGPT4V (Chen et al., 2023), an open-source dataset that integrates data from LLaVA-Instruct (Liu et al., 2024a) and COCO (Lin et al., 2014).
+Originally, ShareGPT4V contained 100,000 captions generated by GPT-4 Vision, later expanding to 1.2 million image-caption pairs using an improved captioning model while maintaining quality and scale.
+To ensure diversity in synthetic data, we process data in batches.
+Each batch consists of 20,000 images and generates 5,000 synthetic conversations, with each conversation containing an average of 4.65 images.
+In total, we use 640,000 image-caption pairs to produce 160,000 synthetic conversations.
+We employ a vector sampling algorithm alongside long-form question prompts to ensure a diverse set of multi-step reasoning tasks.
+The modular design of our pipeline allows easy adaptation to new datasets and applications, making SMiR a scalable solution for multi-image reasoning.
 3 Multi-Image Benchmark
-We create questions that challenge Vision Language Models (VLMs) to reason over multiple related images, analyze relationships, and derive meaning from series of images.
-The diverse range of tasks in this benchmark pushes the boundaries of multimodal understanding, requiring sophisticated analysis and interpretation across various visual contexts.
-Our benchmark compels models to provide detailed explanations for complex visual tasks, going beyond simple multiple-choice answers to demonstrate advanced reasoning capabilities.
-We implement an open-ended questionnaire with a judge model that evaluates responses in a pairwise structure, enabling relative comparisons between models and highlighting subtle differences in their ability to handle complex multi-image reasoning tasks.
+We introduce a benchmark that challenges VLMs to reason over multiple related images, analyze relationships, and derive meaning from image sequences.
+Unlike standard multiple-choice tasks, our benchmark requires detailed explanations, assessing models’ ability to perform complex multimodal reasoning.
+To enable robust comparisons, we employ an open-ended evaluation framework where a judge model ranks responses in a pairwise structure, highlighting subtle differences in reasoning capabilities.
 3.1 Benchmark Overview
-To address these challenges, we developed SMiR-Bench, a multi-turn benchmark consisting of 200 examples across seven diverse topics. These topics include:
+We introduce SMiR-Bench, a multi-turn benchmark consisting of 200 examples spanning seven diverse topics. These topics include:
   • Bird: identifying species and reasoning over distinguishing features
-  • Matching: pairing photos based on visual cues
+  • Matching: pairing photos based on visual similarities
   • Object Character Recognition (OCR): reading and reasoning about academic texts
-  • Pattern: identifying visual task patterns
-  • Ranking: ordering everyday objects based on preferences
-  • Storytelling: narrating based on image sequences, as shown in Figure 4
-  • Visual: finding connections between images
-The benchmark was curated collaboratively by a human annotator and GPT-4, utilizing images from the internet and the Shot2Story dataset (Han et al., 2023).
+  • Pattern: recognizing visual patterns in structured tasks
+  • Ranking: ordering objects based on contextual preferences
+  • Storytelling: narrating events from image sequences, as shown in Fig. 3
+  • Visual: establishing meaningful connections between images
+For each category, we carefully selected 4-5 images from internet sources.
+Storytelling examples were specifically drawn from Shot2Story (Han et al., 2023) videos to leverage their sequential nature.
+GPT-4 was prompted to generate multi-turn question-answer conversations for these images, which were subsequently reviewed and refined by human annotators to ensure coherence and meaningful progression across dialogue turns.
 3.2 Evaluation Methodology
-SMiR-Bench extends the Auto-Hard-Auto v0.1 (Li et al., 2024c) framework to the multimodal domain.
-The evaluation process begins with all models completing the benchmark tasks.
-Subsequently, one model is selected as the baseline, and a GPT-4o judge conducts pairwise comparisons between the baseline model’s responses and those of every other model.
-For each comparison, the judge evaluates responses on helpfulness, relevance, and conciseness.
-The judge’s assessments result in comparative rankings between two models (represented as A and B), such as $A\gg B$ (A significantly better than B), $A>B$ (A better than B), $A=B$ (A equal to B), or their inverses.
-Based on these pairwise comparisons and confidence levels, the framework generates an overall ranking of all models.
-This comprehensive approach enables a multi-turn, automatic, and challenging evaluation process, providing a nuanced assessment of model performance in multimodal tasks.
-By leveraging pairwise comparisons and confidence levels, it offers a clear view of how models perform relative to each other, complete with statistical reliability measures.
+SMiR-Bench extends the Auto-Hard-Auto v0.1 (Li et al., 2024b) framework to multimodal evaluation.
+The process begins with all models completing the benchmark tasks.
+One model is selected as the baseline, and a GPT-4o judge conducts pairwise comparisons between the baseline model’s responses and those of other models.
+The judge evaluates responses based on helpfulness, relevance, and conciseness, assigning comparative rankings such as $A\gg B$ (A significantly better than B), $A>B$ (A better than B), or $A=B$ (A equal to B).
+The framework aggregates these pairwise comparisons and confidence levels to produce an overall ranking of all models.
+An example of this evaluation process is illustrated in Fig.
+3.
+4 Results
+To evaluate the effectiveness of our SMiR dataset, we finetune popular open-source VLMs and assess their performance on SMiR-Bench.
+Our experiments show that models fine-tuned on SMiR achieve up to an 8% improvement over their base versions, demonstrating the dataset’s value for enhancing multi-image reasoning capabilities.
+We begin with two pre-trained models: Mantis-8B-siglip-llama3-pretrained (Jiang et al., 2024) and Idefics-8B (Laurençon et al., 2024b).
+The first model integrates a Llama-3-8B backbone with a SIGLIP encoder and is pre-trained on CC3M (Gan et al., 2022).
+The second employs a Mistral-7B-v0.1 backbone with a perceiver-based projection (Jaegle et al., 2021) and is pre-trained on OBELICS.
+Our evaluation consists of multiple stages.
+First, both pre-trained models undergo instruction tuning on MANTIS-Instruct (721K examples), resulting in Mantis-8B-siglip-llama3 and Mantis-8B-Idefics2.
+We then fine-tune these same pre-trained models on our SMiR dataset, producing SMiR-8B-siglip-llama3 and SMiR-8B-Idefics2.
+To assess model performance, we conduct pairwise comparisons using GPT-4o as a judge model against three baselines:
+  • GPT-4-Turbo, to compare our models against a leading closed-source solution (Table 5).
+  • Mantis-8B-siglip-llama3, to evaluate the improvements from SMiR fine-tuning (Table 3).
+  • Mantis-8B-Idefics2, to assess the performance gains of the Idefics-based model trained on SMiR (Table 4).
+Despite being trained on a significantly smaller dataset, models fine-tuned on SMiR outperform their MANTIS-tuned counterparts across all evaluation scenarios, demonstrating the value of our dataset for improving multi-image reasoning.
+Moreover, the significant performance gap between open-source and closed-source models on SMiR-Bench highlights the benchmark’s difficulty.
+Unlike prior benchmarks, SMiR-Bench focuses on highly correlated images, multi-turn question answering, and open-ended reasoning, making it a more rigorous test of multimodal understanding.
+Table 3: SMiR-Bench scores with Mantis-8B-siglip-llama3 baseline.
+Model Name SMiR-Bench $\Delta$ 95% CI Average Tokens
+Claude-3-Opus-20240229 97.4 (-1.3, 1.0) 321
+Claude-3-5-Sonnet-20240620 97.1 (-1.3, 1.2) 362
+GPT-4-Turbo 96.4 (-1.8, 1.0) 359
+Gemini-1.5-Pro 96.3 (-1.3, 1.3) 361
+GPT-4o 91.5 (-2.4, 1.8) 316
+SMiR-8B-siglip-llama3 58.1 +8.1% (-4.6, 5.3) 156
+Mantis-8B-siglip-llama3 50.0 (0.0, 0.0) 146
+LLaVA-v1.6-Mistral-7B-HF 15.8 (-2.4, 3.1) 317
+Mantis-8B-siglip-llama3-pretraind 8.7 (-2.1, 2.1) 207
+Table 4: SMiR-Bench scores with Mantis-8B-Idefics2 baseline.
+Model Name SMiR-Bench $\Delta$ 95% CI Average Tokens
+Claude-3-Opus-20240229 98.1 (-1.2, 0.9) 321
+Claude-3-5-Sonnet-20240620 98.1 (-1.1, 1.1) 362
+Gemini-1.5-Pro 97.5 (-1.1, 0.9) 361
+GPT-4-Turbo 95.9 (-1.5, 1.6) 359
+GPT-4o 94.5 (-1.6, 1.5) 316
+SMiR-8B-Idefics2 58.0 +8.0% (-5.3, 3.9) 157
+Mantis-8B-Idefics2 50.0 (0.0, 0.0) 171
+Idefics2-8B 29.7 (-3.7, 4.4) 118
+LLaVA-v1.6-Mistral-7B-HF 16.7 (-3.4, 3.0) 317
+5 Conclusion
+We introduces SMiR, a synthetic data pipeline for enhancing multi-image reasoning in open-source VLMs.
+By leveraging multimodal embeddings and grouping algorithms, our approach generates high-quality instruction tuning data, leading to up to an 8% improvement on SMiR-Bench.
+However, certain limitations must be acknowledged.
+Random sampling with iteration is computationally intensive due to the need to recalculate distance embeddings for each newly sampled image.
+The scalability of our synthetic data also requires further investigation.
+Additionally, while SMiR-Bench provides a rigorous evaluation, our approach should be tested across other existing benchmarks to assess its generalizability.
+Appendix A Related Works
+Vision Language Models
+We focus on instruction tuning Vision-Language Models (VLMs) that utilize a pretrained Large Language Model (LLM) backbone because this approach is cost-effective and more accessible for the open-source community.
+Since the backbone responsible for language understanding is already trained, the overall training process becomes simpler and requires fewer resources.
+Our primary task involves aligning the vision encoder—typically architectures like Vision Transformer (ViT) (Dosovitskiy, 2020), SigLIP (Zhai et al., 2023), or CLIP (Radford et al., 2021)—with the LLM backbone.
+This alignment is facilitated through linear layers that connect the vision encoder to the backbone, enabling the integration of visual and textual information.
+For instance, BLIP-2 (Li et al., 2023b) uses OPT (Zhang et al., 2022) and FLAN-T5 (Chung et al., 2022) as backbones, MiniGPT-4 (Zhu et al., 2023) utilizes Vicuna (Chiang et al., 2023), and Qwen-2-VL (Wang et al., 2024) employs Qwen-2-1.5B (Yang et al., 2024) as the language backbone.
+In this paper, we focus on creating a high-quality multi-image reasoning dataset for instruction tuning instead of large-scale interleaved pretraining datasets like OBELICS (Laurençon et al., 2024a), MINT-1T (Awadalla et al., 2024), and LAION-5B (Schuhmann et al., 2022).
+Multi-Image Reasoning Data
+Recent advancements in multi-image reasoning instruction tuning datasets include MANTIS (Jiang et al., 2024) and MMDU-45K (Liu et al., 2024f), both aiming to improve reasoning capabilities in VLMs.
+However, these datasets have limitations in their approaches.
+MANTIS randomly concatenates single image pairs from LLaVA-665k (Liu et al., 2024a), which often results in uncorrelated images within multi-image scenarios, potentially undermining the complexity of reasoning tasks.
+MMDU-45K attempts to address this issue by utilizing sentence transformers (Reimers, 2019) with description text and clustering techniques to group related images, but does not consider visual components.
+The dataset is then further enhanced, assisted by GPT-4 to generate comprehensive answers for the grouped images.
+Building upon these efforts, SMiR introduces a novel approach that leverages both vital visual and caption information to ensure highly correlated images within multi-image sets with the use of open-source LLMs.
+These scalable methods leads to the generation of more challenging questions that require deeper analysis and understanding of visual relationships, pushing the boundaries of multi-image reasoning capabilities in VLMs.
+Multi-Image Reasoning Benchmarks
+Recent VLM benchmarks (Chiang et al., 2024; Lin et al., 2024; Liu et al., 2024d) have made strides by incorporating free-response evaluations, marking a significant improvement over traditional multiple-choice formats.
+However, these benchmarks still lack a comprehensive approach that combines automatic, multi-turn, and pairwise evaluation capabilities.
+Our benchmark addresses this gap, drawing inspiration from Auto-Hard-Auto v0.1 (Li et al., 2024b).
+We have adapted and expanded this framework to enable robust multimodal evaluation, providing a more holistic assessment of VLM performance across complex, multi-image reasoning tasks.
+This approach allows for a deeper analysis of both the final answers and the underlying reasoning processes employed by VLMs in real-world SMiR-Bench scenarios.
 Figure 3: Evaluation Benchmark (Storytelling) Using GPT-4o as Judge
-4.3 Results
-To evaluate the effectiveness of our SMiR dataset, we finetune popular open-source VLMs and assess their performance on SMiR-Bench.
-Our experiments demonstrate significant improvements, with finetuned models achieving up to 8% better performance compared to their base versions.
-4.4 Experiments
-For our evaluation, we begin with two pre-trained models: Mantis-8B-siglip-llama3-pretrained (Jiang et al., 2024) and idefics-8b (Laurençon et al., 2024b).
-The first model combines a Llama-3-8B backbone with a SIGLIP encoder and is pretrained on CC3M (Gan et al., 2022).
-The second uses a Mistral-7B-v0.1 backbone with a perceiver (Jaegle et al., 2021) and MLP projection, pretrained on OBELICS.
-Both models were initially instruction-tuned on MANTIS-Instruct (721K examples) to create Mantis-8B-siglip-llama3 and Mantis-8B-Idefics2.
-We then finetune these same pre-trained models on our SMiR dataset to create SMiR-8B-siglip-llama3-160 (trained on 160K examples) and SMiR-8B-Idefics2-100 (trained on 100K examples).
-To evaluate performance, we conduct pairwise comparisons using GPT-4 as a judge model against three specific baselines: GPT-4-Turbo to assess our models against a closed-source solution (Table 3), Mantis-8B-siglip-llama3 to evaluate SMiR-8B-siglip-llama3-160 (Table 4), and Mantis-8B-Idefics2 to compare with SMiR-8B-Idefics2-100 (Table 5).
-Remarkably, our SMiR-tuned models outperform their MANTIS-tuned counterparts despite being trained on a significantly smaller dataset.
-These results demonstrate the effectiveness of our SMiR dataset for finetuning, showing consistent performance improvements across all evaluation scenarios.
-Table 3: SMiR-Bench Scores with GPT-4-Turbo Baseline
+Table 5: SMiR-Bench scores with GPT-4-Turbo baseline.
 Model Name SMiR-Bench $\Delta$ 95% CI Average Tokens
@@
 Claude-3-Opus-20240229 34.4 (-4.6, 4.4) 321
-SMiR-8B-siglip-llama3-160 5.0 +.8% (-2.2, 2.7) 183
+SMiR-8B-siglip-llama3 5.0 +.8% (-2.2, 2.7) 183
 Mantis-8B-siglip-llama3 4.2 (-1.2, 2.2) 146
-SMiR-8B-Idefics2-100 4.5 +1.3% (-1.3, 1.4) 157
+SMiR-8B-Idefics2 4.5 +1.3% (-1.3, 1.4) 157
 Mantis-8B-Idefics2 3.2 (-1.2, 1.2) 171
@@
 LLaVA-v1.6-mistral-7b-hf 1.1 (-.7, .7) 317
-Table 4: SMiR-Bench Scores with Mantis-8B-siglip-llama3 Baseline
-Model Name SMiR-Bench $\Delta$ 95% CI Average Tokens
-Claude-3-Opus-20240229 97.4 (-1.3, 1.0) 321
-Claude-3-5-Sonnet-20240620 97.1 (-1.3, 1.2) 362
-GPT-4-Turbo 96.4 (-1.8, 1.0) 359
-Gemini-1.5-Pro 96.3 (-1.3, 1.3) 361
-GPT-4o 91.5 (-2.4, 1.8) 316
-SMiR-8B-siglip-llama3-160 58.1 +8.1% (-4.6, 5.3) 156
-Mantis-8B-siglip-llama3 50.0 (0.0, 0.0) 146
-LLaVA-v1.6-Mistral-7B-HF 15.8 (-2.4, 3.1) 317
-Mantis-8B-siglip-llama3-pretraind 8.7 (-2.1, 2.1) 207
-Table 5: SMiR-Bench Scores with Mantis-8B-Idefics2 Baseline
-Model Name SMiR-Bench $\Delta$ 95% CI Average Tokens
-Claude-3-Opus-20240229 98.1 (-1.2, 0.9) 321
-Claude-3-5-Sonnet-20240620 98.1 (-1.1, 1.1) 362
-Gemini-1.5-Pro 97.5 (-1.1, 0.9) 361
-GPT-4-Turbo 95.9 (-1.5, 1.6) 359
-GPT-4o 94.5 (-1.6, 1.5) 316
-SMiR-8B-Idefics2-100 58.0 +8.0% (-5.3, 3.9) 157
-Mantis-8B-Idefics2 50.0 (0.0, 0.0) 171
-Idefics2-8B 29.7 (-3.7, 4.4) 118
-LLaVA-v1.6-Mistral-7B-HF 16.7 (-3.4, 3.0) 317
-Figure 5: SMiR Dataset Size vs. Benchmark Score
-5 Conclusion
-This paper introduces a synthetic data pipeline designed to enhance multi-image reasoning capabilities on open-source VLMs.
-By leveraging multimodal embeddings and grouping algorithms, the pipeline generates high-quality synthetic multi-image reasoning instruction tuning data.
-The approach yields up to 8% improvement on SMiR-Bench for popular open-source models, demonstrating the significant potential of synthetic data in advancing open-source VLM models.
-Limitations
-Our methods have several limitations.
-Random sampling with iteration is time-intensive due to the need for recalculating distance embeddings for each new image sampled.
-Further investigation is needed to determine the scalability of our synthetic data.
-Future research should focus on developing more time-efficient algorithms and optimizing data mixtures.
-Broader Impact
-This paper introduces a method for generating high-quality, cost-effective data for VLMs, addressing the growing challenge of data scarcity.
-By advancing these open-source techniques, we contribute to narrowing the performance gap between open and closed-source models, promoting more accessible and powerful multimodal AI.
-Appendix A Algorithm Details
-A.1 Greedy Cluster Matching Algorithm
+Appendix B Algorithm Details
+B.1 Greedy Cluster Matching Algorithm
 We present the pseudocode for the Greedy Cluster Matching in Algorithm 1.
@@
 […]
-A.2 Random Sampling with Iteration
+B.2 Random Sampling with Iteration
+Figure 4: Sampling based on distances between multimodal embeddings of image-caption pairs.
 We present the pseudocode for the Random Sampling with Iteration in Algorithm 2.
@@
 26:return $selected$
-Appendix B Data Samples
+Appendix C Data Samples
 For the sake of brevity, we have included only two examples from the multi-image data we have generated, featuring questions with a related theme from both the Greedy Cluster Matching Algorithm and Random Sampling with Iteration, highlighting how the algorithms differ in their approach to generating questions.
-B.1 Greedy Cluster Matching Algorithm
+C.1 Greedy Cluster Matching Algorithm
+We present two novel algorithms, Greedy Cluster Matching and Random Sampling with Iteration, designed to group correlated images prior to leveraging an open-source Large Language Model (LLM) for synthetic data generation in multi-image reasoning tasks.
+The emphasis on correlated images is crucial, as it facilitates challenging multi-image reasoning scenarios.
+These scenarios require the model to identify intricate relationships and differentiate between visually similar scenes, thus enhancing the complexity and realism of the reasoning process.
+Greedy Cluster Matching
+After employing HDBSCAN (Malzer & Baum, 2020), a density-based clustering algorithm, to group the SigLIP and CLIP multimodal embeddings into coherent clusters, we developed a greedy algorithm to establish meaningful relationships between the two embedding spaces.
+Given cluster sets $C_{S}=\{S_{1},...,S_{m}\}$ from SigLIP and $C_{C}=\{C_{1},...,C_{n}\}$ from CLIP embeddings, the algorithm:
+1) Orders clusters by size in descending order
+2) Iteratively selects the largest remaining cluster $X_{max}=\arg\max_{X\in C_{S}\cup C_{C}}|X|$
+3) For the selected cluster, finds its best match $Y_{best}$ from the other embedding space using: $Y_{best}=\arg\max_{Y}\frac{|X_{max}\cap Y|}{\frac{|X_{max}|+|Y|}{2}}$
+This process continues until all clusters are matched or one set is exhausted, demonstrating high efficacy in producing quality image-caption pairs through dual-embedding validation.
+The score function, which measures overlap normalized by average cluster size, ensures each cluster corresponds to a semantically similar cluster from the other model.
+The detailed steps of this algorithm and pseudocode are presented in Algorithm 1 (Section B.1).
+While this approach effectively leverages both embedding models to confirm spatial relationships and associated semantic meanings, with images matched within clusters being corroborated by two independent embedding models, it has a notable limitation: sampling confined to matched clusters can lead to overly specialized image subjects, such as clustering only sheep-related images rather than diverse animal scenes, as demonstrated in Fig.
+5 (Section C.1).
+This specialization occurs because sampling is restricted to a single matched cluster, potentially limiting the diversity of selected images.
 Samples obtained through Greedy Cluster Matching typically feature similar subjects and shot compositions, but when paired with carefully crafted prompts, these similarities can be leveraged to generate more challenging and nuanced questions.
 Figure 5: Images sampled from the same matched cluster often feature similar subjects or scenes.
-B.2 Random Sampling with Iteration
+C.2 Random Sampling with Iteration
 Random sampling tends to yield greater diversity in subjects compared to Greedy Clustering, striking a balance between variety and relatedness that can potentially lead to more robust and wide-ranging question sets.
 Figure 6: Images sampled using the iterative algorithm allow for different yet related subjects (e.g., various animal species)
-Appendix C Prompt
+Appendix D Prompt
 While prompts play a crucial role in data generation, optimizing them remains a significant challenge.
 After numerous iterations, we identified two particularly effective prompts for multi-image data generation.
-C.1 LLaVA Style Prompt
+D.1 LLaVA Style Prompt
 Inspired by LLaVA (Liu et al., 2024b), our approach utilizes a specialized prompt to address simpler multi-image and single-image tasks, focusing on more straightforward visual comprehension and analysis.
@@
 Return your challenging and complex question and follow-up questions in the format: ‘‘User: [INSERT QUESTION], Assistant: [INSERT ANSWER] User: [INSERT FOLLOW-UP QUESTION 1], Assistant: [INSERT FOLLOW-UP ANSWER 1] User: [INSERT FOLLOW-UP QUESTION 2], Assistant: [INSERT FOLLOW-UP ANSWER 2] User: [INSERT FOLLOW-UP QUESTION 3], Assistant: [INSERT FOLLOW-UP ANSWER 3]...’’
-C.2 Longer Prompt
+D.2 Longer Prompt
 Our approach aims to generate more complex, multi-turn questions that require in-depth reasoning across multiple images.